# Probing quantum entanglement with Generalized Parton Distributions at the Electron-Ion Collider

### v2

@@ IV.1 Polarization
First we show the results on the polarization of the quark $B_{n}$ and the antiquark $\bar{B}_{n}$. As we noted already, $B_{n}$ and $\bar{B}_{n}$ are nonvanishing only for massive quarks and for the transversely polarized photon, and tend to be suppressed at high CM energies. Therefore, it is best to focus on low-$W$ UPC or low-$W$ DIS at low-$Q^{2}$. In Fig. 3, we plot $B_{n}$ and $\bar{B}_{n}$ for the strange (upper panels) and charm (lower panels) quarks in UPC at $W=30$ GeV. We find that $B_{n}$ is negative and $\bar{B}_{n}$ is positive for all massive quark flavors $q=s,c,b$, consistently with (41). The net polarization along the $\hat{n}$ direction $B_{n}(z)+\bar{B}_{n}(z)\propto z-\bar{z}$ vanishes at $z=\frac{1}{2}$. The peak structure in $k_{\perp}$ is as expected from the analytic formula (34), with the peak magnitude reaching 80%, 60%, 50% for strange, charm and bottom quarks, respectively. [^5] It should be mentioned that $B_{n}$ of the strange quark is largest in the low-momentum region $k_{\perp}\gtrsim 1$ GeV where, strictly speaking, the present perturbative approach is not fully justified in photoproduction $Q=0$. These peaks occur when the quark (antiquark) is ‘soft’ $z\ll 1$ ($\bar{z}\ll 1$) in the original frame (2) in which the photon is fast moving and the $q\bar{q}$ pair is produced in the forward rapidity region, cf., the comment below (35). Negative $B_{n}$ means that the quark is polarized in the direction of $\vec{q}\times\vec{k}_{q}$, see Fig. 1. Similarly, the antiquark is polarized along the direction $\vec{q}\times\vec{k}_{\bar{q}}$.
@@
Comments are in order regarding the magnitude of the polarization. In semi-inclusive quark or antiquark production, the imaginary part of the amplitude responsible for nonvanishing polarization is suppressed by a factor of $\alpha_{s}$ since it requires a loop. Typically, one finds (sub-)percent level asymmetries in such perturbative calculations Dharmaratna:1996xd; Benic:2024fvk. However, in exclusive production with color singlet exchanges in the $t$-channel, the real and imaginary parts of the GPD amplitudes are of the same order in $\alpha_{s}$. Consequently, $B_{n},\bar{B}_{n}$ (34) are order unity instead of order $\alpha_{s}$. Still, a polarization as large as 50-80% is quite remarkable. To our knowledge, such a strong polarization of perturbative origin has not been reported in the QCD literature.
@@
It is generally expected that the quark-level polarization is largely retained during fragmentation into heavy baryons $q\to\Lambda_{q}$, $\bar{q}\to\bar{\Lambda}_{q}$ Galanti:2015pqa. In particular, the process $s\to\Lambda\to p+\pi^{-}$ can be used as a polarimeter of the strange quark. The HERMES collaboration HERMES:2007fpi measured the polarization of $\Lambda$ and $\bar{\Lambda}$ in semi-inclusive DIS in the photoproduction region $Q\approx 0$. They found a positive polarization for $\Lambda$ but almost zero polarization for $\bar{\Lambda}$. On the other hand, the COMPASS collaboration COMPASS:2021bws found very small values consistent with zero for both $\Lambda$ and $\bar{\Lambda}$. The contribution from exclusive $s\bar{s}$ production, if any at all in these low energy experiments, will be significantly diluted by the inclusive background. On the other hand, diffractive events with a rapidity gap are copious at the EIC. By triggering on such events, one should be able to test our prediction in future.
@@
IV.2 Entanglement
The degree of entanglement between the quark and antiquark spins can be quantified by $\Delta_{1,2}$ introduced in (51). First, consider UPCs at the LHC and fix $W=500$ GeV, $Q=0$ and $k_{\perp}^{\rm max}=30$ GeV. In Fig. 4, we plot the left hand side of (50) in the $(z,k_{\perp})$ plane for $u\bar{u}$ (left), $c\bar{c}$ (middle) and $b\bar{b}$ (right) pairs, respectively. The results for $d\bar{d}$, $s\bar{s}$ pairs are similar to the $u\bar{u}$ case. In practice, these are plots of $\Delta_{2}^{T}$ that dominates in (50). Since $W$ is typically quite large at the LHC, the results are close to those in the Regge limit calculation in Fucilla:2025kit. In particular, the pairs are everywhere entangled (i.e., (50) is always satisfied). Light quarks are maximally entangled ($\Delta^{T}_{2}\approx 2$) along the line $z=\frac{1}{2}$ as discussed around (58). For massive quarks, an almost maximally entangled state
@@ IV.2 Entanglement
is realized around $z=\frac{1}{2}$ and $k_{\perp}\approx m$. As explained in Fucilla:2025kit, this occurs because $I_{g}^{L}\approx 0$ (see the comment below (46)) and hence $C_{nn}^{T}\approx-1$ in this region. The peak is followed by a dip as $k_{\perp}$ increases (i.e., entanglement becomes weak) before rising again at higher $k_{\perp}$ where the mass effect becomes less important.
Next, we consider DIS at $W=30$ GeV and $Q=10$ GeV. In Fig. 5, we plot the left hand side of (50) (with $T\to{\rm DIS}$ in the superscript) for $u\bar{u}$ (left), $s\bar{s}$ (middle), $b\bar{b}$ (right) pairs. The $u\bar{u}$ plot is asymmetric under $z\leftrightarrow\bar{z}$. This is because the valence quark integrals $I_{L}^{q}$, $I_{T}^{q_{1,2}}$ with $q=u,d$ do not have this symmetry. Actually, the same is true in Fig. 4 (left), but the asymmetry is barely noticeable because at high energy the cross section is dominated by the gluon contribution which is symmetric. Since $Q$ is relatively high, the results reflect a significant interference between the transverse and longitudinal photon contributions. In $b\bar{b}$ (and $c\bar{c}$) production, we find narrow white regions where the pair is not entangled (i.e., the pair is separable). We have studied other sets of $(W,Q)$ and found that regions of no-entanglement appear only for $c\bar{c},b\bar{b}$ pairs in small patches at most. This means that, even though the conditions (53) do not hold away from the Regge limit, entanglement remains a robust feature of exclusive $q\bar{q}$ production.
@@
Figure 4: Left hand side of (50) in UPC at $W=500$ GeV, $Q=0$. Left: $u\bar{u}$ pair, Middle: $c\bar{c}$ pair, Right: $b\bar{b}$ pair. The pairs are almost maximally entangled $\Delta_{2}^{T}\approx 2$ in the red region.
@@
Figure 5: Left hand side (50) in DIS at $Q=10$ GeV, $W=30$ GeV. Left: $u\bar{u}$ pair, Middle: $s\bar{s}$ pair, Right: $b\bar{b}$ pair. The pair is not entangled in the white region.
IV.3 Bell nonlocality
Next, we delineate the regions where the $q\bar{q}$ pair exhibits Bell nonlocality. In Fig. 6 (left), we plot $\mu_{1}+\mu_{2}-1$ (see (60)) for $b\bar{b}$ pairs in the same UPC kinematics as in Fig. 4 (right). We see a large white region where $\mu_{1}+\mu_{2}-1<0$, meaning that the Bell-CHSH inequality is not violated there. This is in contrast to the finding in Fucilla:2025kit that, in the Regge limit, the pair exhibits Bell nonlocality in the entire kinematical region of $(z,k_{\perp})$. It is also in contrast to Fig. 4 (right) which shows that the pair remains entangled even away from the Regge limit. These observations are consistent with the fact that, in general, Bell nonlocality is a sufficient but not necessary condition for entanglement.
In the middle and right panels of Fig. 6, we plot the same quantity for the $s\bar{s}$ and $b\bar{b}$ pairs in DIS ($W=30$ GeV, $Q=10$ GeV) to be directly compared to the middle and right panels in Fig. 5. Again we find regions where Bell nonlocality is lost, even though the pairs are still entangled.
@@ IV.3 Bell nonlocality
Figure 6: Plot of $\mu_{1}+\mu_{2}-1$, see (60). The Bell-CHSH inequality is not violated in the white regions. Left: $b\bar{b}$ production in UPC at $W=500$ GeV, $Q=0$, compare with Fig. 4 (right). Middle: $s\bar{s}$ production in DIS, compare with Fig. 5 (middle). Right: $b\bar{b}$ pair production in DIS, compare with Fig. 5 (right).
IV.4 Magic
Finally, in Fig. 7 and Fig. 8, we plot the stabilizer Rényi entropy $M_{2}$ (64) in UPC and DIS, respectively, for different quark flavors. As expected, $M_{2}$ is suppressed when entanglement is nearly maximal. But other than this general statement, it is difficult to anticipate in which kinematical regions $M_{2}$ is enhanced. We however see that $M_{2}$ tends to be larger for heavier quarks, with the maximal value around $0.58$ for the $b\bar{b}$ pair in DIS with $W=30$ GeV, $Q=3$ GeV, see Fig. 8 (right). Curiously, in the entire parameter space we have explored, $M_{2}$ is always smaller than (70), although in principle it can exceed this bound because $B_{n},\bar{B}_{n}\neq 0$.
@@ IV.4 Magic
Figure 8: Stabilizer Rényi entropy (64) in DIS ($W=30$ GeV, $Q=10$ GeV) for $s\bar{s}$ (left), $c\bar{c}$ (middle), and $b\bar{b}$ (right) pairs.
V Conclusions
In this paper, we have analyzed exclusive quark-antiquark production in DIS and UPCs from the viewpoint of Quantum Information Science (QIS). This is a generalization of the previous work in the high energy (Regge) limit Fucilla:2025kit to more practical CM energies at the EIC where collinear factorization with GPDs is an appropriate theoretical tool. The scattering amplitudes computed in the GPD framework are complex, in contrast to being purely real in the one-gluon exchange approximation Qi:2025onf and purely imaginary in the Regge limit. An immediate consequence is that massive $q\bar{q}$ pairs are transversely polarized without the usual $\alpha_{s}$ suppression in inclusive production. In low-energy UPC or DIS, the polarization can easily exceed 50% in certain kinematic regions. Moreover, we have observed rich patterns of entanglement, Bell nonlocality and magic due to the interference between the real and imaginary parts of the GPD amplitudes, and also between the transverse and longitudinal photons. Together with the EIC’s ability to vary $W$ and $Q$ over an unprecedentedly wide kinematical range, this makes DIS a particularly fascinating arena for exploring QIS in collider physics.
We have focused on the polarization and entanglement of the produced $q\bar{q}$ pairs. Whether these properties survive as observables in realistic experiments is a nontrivial problem that requires separate theoretical investigations. For massive quarks $s,c,b$ (or antiquarks $\bar{s},\bar{c},\bar{b}$), the canonical method is to tag events where the quarks fragment into heavy baryons $\Lambda,\Lambda_{c},\Lambda_{b}$ (or antibaryons $\bar{\Lambda},\bar{\Lambda}_{c},\bar{\Lambda}_{b}$) Galanti:2015pqa. Their subsequent weak decays such as $\Lambda\to p+\pi^{-}$ allows one to reconstruct the polarization vectors $B_{a},\bar{B}_{a}$ Bunce:1976yb; Lundberg:1989hw and the spin density matrix $C_{ab}$ Tornqvist:1980af; Afik:2025grr. For light quarks, methods outlined in Cheng:2025cuv may be applicable to DIS. In both cases, detailed simulations with a realistic detector setup at the EIC are necessary Lin:2025eci; Gu:2025ijz.
@@ Appendix A Maximal magic without polarization
In this Appendix, we determine the upper limit of the stabilizer Rényi entropy Leone:2021rzd for generic two qubit systems with vanishing polarization. Following Liu:2025frx, we parametrize pure two-qubit states as
@@
It is easy to check that when (82) is satisfied, (80) is automatically satisfied. The maximal value of $M_{2}$ for systems with $\vec{B}_{1}=\vec{B}_{2}=0$ is therefore obtained by minimizing (16) of Liu:2025frx with the constraints (82). There are many degenerate minima. For example, we find
@@
Minimizing (16) of Liu:2025frx under this additional constraint, we find the extremal condition $\cos(4\phi_{1})=-1$, or equivalently,

### v3

@@ V.1 Polarization
First we show the results on the polarization of the quark $B_{n}$ and the antiquark $\bar{B}_{n}$. As we noted already, $B_{n}$ and $\bar{B}_{n}$ are nonvanishing only for massive quarks and for the transversely polarized photon, and tend to be suppressed at high CM energies. Therefore, it is best to focus on low-$W$ UPC or low-$W$ electroproduction at low-$Q^{2}$. In Fig. 3, we plot $B_{n}$ and $\bar{B}_{n}$ for the strange (upper panels) and charm (lower panels) quarks in UPC at $W=30$ GeV. We find that $B_{n}$ is negative and $\bar{B}_{n}$ is positive for all massive quark flavors $q=s,c,b$, consistently with (46). The net polarization along the $\hat{n}$ direction $B_{n}(z)+\bar{B}_{n}(z)\propto z-\bar{z}$ vanishes at $z=\frac{1}{2}$. The peak structure in $k_{\perp}$ is as expected from the analytic formula (39), with the peak magnitude reaching 80%, 60%, 50% for strange, charm and bottom quarks, respectively. [^6] It should be mentioned that $B_{n}$ of the strange quark is largest in the low-momentum region $k_{\perp}\gtrsim 1$ GeV where, strictly speaking, the present perturbative approach is not fully justified in photoproduction $Q=0$. These peaks occur when the quark (antiquark) is ‘soft’ $z\ll 1$ ($\bar{z}\ll 1$) in the original frame (2) in which the photon is fast moving and the $q\bar{q}$ pair is produced in the forward rapidity region, cf., the comment below (40). Negative $B_{n}$ means that the quark is polarized in the direction of $\vec{q}\times\vec{k}_{q}$, see Fig. 1. Similarly, the antiquark is polarized along the direction $\vec{q}\times\vec{k}_{\bar{q}}$.
@@
Comments are in order regarding the magnitude of the polarization. In semi-inclusive quark or antiquark production, the imaginary part of the amplitude responsible for nonvanishing polarization is suppressed by a factor of $\alpha_{s}$ since it requires a loop. Typically, one finds (sub-)percent level asymmetries in such perturbative calculations Dharmaratna and Goldstein (1996); Benić et al. (2024). However, in exclusive production with color singlet exchanges in the $t$-channel, the real and imaginary parts of the scattering amplitudes are of the same order in $\alpha_{s}$. Consequently, $B_{n},\bar{B}_{n}$ (39) are order unity instead of order $\alpha_{s}$. Still, a polarization as large as 50-80% is quite remarkable. To our knowledge, such a strong polarization of perturbative origin has not been reported in the QCD literature.
@@
It is generally expected that the quark-level polarization is largely retained during fragmentation into heavy baryons $q\to\Lambda_{q}$, $\bar{q}\to\bar{\Lambda}_{q}$ Galanti et al. (2015). In particular, the process $s\to\Lambda\to p+\pi^{-}$ can be used as a polarimeter of the strange quark. The HERMES collaboration Airapetian et al. (2007) measured the polarization of $\Lambda$ and $\bar{\Lambda}$ in semi-inclusive DIS in the photoproduction region $Q\approx 0$. They found a positive polarization for $\Lambda$ but almost zero polarization for $\bar{\Lambda}$. On the other hand, the COMPASS collaboration Alexeev et al. (2022) found very small values consistent with zero for both $\Lambda$ and $\bar{\Lambda}$. The contribution from exclusive $s\bar{s}$ production, if any at all in these low energy experiments, will be significantly diluted by the inclusive background. On the other hand, diffractive events with a rapidity gap are copious at the EIC. By triggering on such events, one should be able to test our prediction in future.
@@
V.2 Entanglement
The degree of entanglement between the quark and antiquark spins can be quantified by $\Delta_{1,2}$ introduced in (56). First, consider UPCs at the LHC and fix $W=500$ GeV, $Q=0$ and $k_{\perp}^{\rm max}=30$ GeV. In Fig. 4, we plot the left hand side of (55) in the $(z,k_{\perp})$ plane for $u\bar{u}$ (left), $c\bar{c}$ (middle) and $b\bar{b}$ (right) pairs, respectively. The results for $d\bar{d}$, $s\bar{s}$ pairs are similar to the $u\bar{u}$ case. In practice, these are plots of $\Delta_{2}^{T}$ that dominates in (55). Since $W$ is typically quite large at the LHC, the results are close to those in the Regge limit calculation in Fucilla and Hatta (2026). In particular, the pairs are everywhere entangled (i.e., (55) is always satisfied). Light quarks are maximally entangled ($\Delta^{T}_{2}\approx 2$) along the line $z=\frac{1}{2}$ as discussed around (63). For massive quarks, an almost maximally entangled state
@@ V.2 Entanglement
is realized at a peak around $z=\frac{1}{2}$ and $k_{\perp}\approx m$. The peak occurs because $I_{g}^{L}\approx 0$ (see the comment below (51)) and hence $C_{nn}^{T}\approx-1$ in this region. The peak is followed by a dip as $k_{\perp}$ increases (i.e., entanglement becomes weak) before rising again at higher $k_{\perp}$ where the mass effect becomes less important. It should be noted that, when $W=500$ GeV and $k_{\perp}$ is a few GeV, typically the region $x\sim\xi\sim 10^{-5}$ of the proton wavefunction is probed. At such small values of $x$, one may have to switch to the Regge limit calculation Fucilla and Hatta (2026) including the gluon saturation Gelis et al. (2010). The two calculations are qualitatively similar, though. In particular, the peak structure around $k_{\perp}\sim m$ and $z\sim\frac{1}{2}$ was also found in Fucilla and Hatta (2026). It would be interesting to perform a detailed comparison of the spin density matrix with or without gluon saturation and small-$x$ evolution effects.
Next, we consider electroproduction at $W=30$ GeV and $Q=10$ GeV. In Fig. 5, we plot the left hand side of (55) (with $T\to{\rm DIS}$ in the superscript) for $u\bar{u}$ (left), $s\bar{s}$ (middle), $b\bar{b}$ (right) pairs. The $u\bar{u}$ plot is asymmetric under $z\leftrightarrow\bar{z}$. This is because the valence quark integrals $I_{L}^{q}$, $I_{T}^{q_{1,2}}$ with $q=u,d$ do not have this symmetry. Actually, the same is true in Fig. 4 (left), but the asymmetry is barely noticeable because at high energy the cross section is dominated by the gluon contribution which is symmetric. Since $Q$ is relatively high, the results reflect a significant interference between the transverse and longitudinal photon contributions. In $b\bar{b}$ (and $c\bar{c}$) production, we find narrow white regions where the pair is not entangled (i.e., the pair is separable). We have studied other sets of $(W,Q)$ and found that regions of no-entanglement appear only for $c\bar{c},b\bar{b}$ pairs in small patches at most. This means that, even though the conditions (58) do not hold away from the Regge limit, entanglement remains a robust feature of exclusive $q\bar{q}$ production.
@@
Figure 4: Left hand side of (55) in UPC at $W=500$ GeV, $Q=0$. Left: $u\bar{u}$ pair, Middle: $c\bar{c}$ pair, Right: $b\bar{b}$ pair. Entanglement is stronger in brighter regions where mostly ${\rm max}\{\Delta_{1},\Delta_{2},|C_{nn}-1\}=\Delta_{2}$. The pair reaches maximally entangled states $\Phi^{-}$ (62) and $\Psi_{n}^{-}$ (80) with $\Delta_{2}\approx 2$ in the indicated regions.
@@
Figure 5: Left hand side (55) in electroproduction at $Q=10$ GeV, $W=30$ GeV. Left: $u\bar{u}$ pair, Middle: $s\bar{s}$ pair, Right: $b\bar{b}$ pair. The pair is not entangled in the white region.
V.3 Bell nonlocality
Next, we delineate the regions where the $q\bar{q}$ pair exhibits Bell nonlocality. In Fig. 6 (left), we plot $\mu_{1}+\mu_{2}-1$ (see (65)) for $b\bar{b}$ pairs in the same UPC kinematics as in Fig. 4 (right). We see a large white region where $\mu_{1}+\mu_{2}-1<0$, meaning that the Bell-CHSH inequality is not violated there. This is in contrast to the finding in Fucilla and Hatta (2026) that, in the Regge limit, the pair exhibits Bell nonlocality in the entire kinematical region of $(z,k_{\perp})$. It is also in contrast to Fig. 4 (right) which shows that the pair remains entangled even away from the Regge limit. These observations are consistent with the fact that, in general, Bell nonlocality is a sufficient but not necessary condition for entanglement. Since the violation of the Bell-CHSH inequality requires stronger quantum correlations than entanglement, it can be seen in a narrower region of phase space.
In the middle and right panels of Fig. 6, we plot the same quantity for the $s\bar{s}$ and $b\bar{b}$ pairs in electroproduction ($W=30$ GeV, $Q=10$ GeV) to be directly compared to the middle and right panels in Fig. 5. Again we find regions where Bell nonlocality is lost, even though the pairs are still entangled.
@@ V.3 Bell nonlocality
Figure 6: Plot of $\mu_{1}+\mu_{2}-1$, see (65). The Bell-CHSH inequality is not violated in the white regions. Left: $b\bar{b}$ production in UPC at $W=500$ GeV, $Q=0$, compare with Fig. 4 (right). Middle: $s\bar{s}$ production in electroproduction, compare with Fig. 5 (middle). Right: $b\bar{b}$ pair production in electroproduction, compare with Fig. 5 (right).
V.4 Magic
Finally, in Fig. 7 and Fig. 8, we plot the stabilizer Rényi entropy $M_{2}$ (69) in UPC and electroproduction, respectively, for different quark flavors. As expected, $M_{2}$ is suppressed when entanglement is nearly maximal. But other than this general statement, it is difficult to anticipate in which kinematical regions $M_{2}$ is enhanced. We however see that $M_{2}$ tends to be larger for heavier quarks, with the maximal value around $0.58$ for the $b\bar{b}$ pair in electroproduction with $W=30$ GeV, $Q=3$ GeV, see Fig. 8 (right). Curiously, in the entire parameter space we have explored, $M_{2}$ is always smaller than (75), although in principle it can exceed this bound because $B_{n},\bar{B}_{n}\neq 0$.
@@ V.4 Magic
Figure 8: Stabilizer Rényi entropy (69) in electroproduction ($W=30$ GeV, $Q=10$ GeV) for $s\bar{s}$ (left), $c\bar{c}$ (middle), and $b\bar{b}$ (right) pairs.
VI Conclusions
In this paper, we have analyzed exclusive quark-antiquark production in electroproduction and UPCs from the viewpoint of Quantum Information Science (QIS). This is a generalization of the previous work in the high energy (Regge) limit Fucilla and Hatta (2026) to more practical CM energies at the EIC where collinear factorization with GPDs is an appropriate theoretical tool. The scattering amplitudes computed in the collinear factorization framework are complex, in contrast to being purely real in the one-gluon exchange approximation Qi et al. (2025) and purely imaginary in the Regge limit. An immediate consequence is that massive $q\bar{q}$ pairs are transversely polarized without the usual $\alpha_{s}$ suppression in inclusive production. In low-energy UPC or electroproduction, the polarization can easily exceed 50% in certain kinematic regions. Moreover, we have observed rich patterns of entanglement, Bell nonlocality and magic due to the interference between the real and imaginary parts of the scattering amplitudes, and also between the transverse and longitudinal photons. Together with the EIC’s ability to vary $W$ and $Q$ over an unprecedentedly wide kinematical range, this makes DIS a particularly fascinating arena for exploring QIS in collider physics.
We have focused on the polarization and entanglement of the produced $q\bar{q}$ pairs to leading order in perturbation theory. In future, the impact of the next-to-leading order corrections should be investigated (cf. Boussarie et al. (2016)). Moreover, other types of GPDs such as the polarized GPDs $\tilde{H}_{q,g}$ can contribute and become important at low energy. Most importantly, whether the entanglement properties discussed in this paper survive as observables in realistic experiments is a nontrivial problem that requires separate theoretical investigations. For massive quarks $s,c,b$ (or antiquarks $\bar{s},\bar{c},\bar{b}$), the canonical method is to tag events where the quarks fragment into heavy baryons $\Lambda,\Lambda_{c},\Lambda_{b}$ (or antibaryons $\bar{\Lambda},\bar{\Lambda}_{c},\bar{\Lambda}_{b}$) Galanti et al. (2015). Their subsequent weak decays such as $\Lambda\to p+\pi^{-}$ allows one to reconstruct the polarization vectors $B_{a},\bar{B}_{a}$ Bunce et al. (1976); Lundberg et al. (1989) and the spin density matrix $C_{ab}$ Tornqvist (1981); Afik et al. (2025b). For light quarks, methods outlined in Cheng and Yan (2025) may be applicable to DIS. In both cases, detailed simulations with a realistic detector setup at the EIC are necessary Lin et al. (2025); Gu et al. (2025).
@@ Appendix A Maximal magic without polarization
In this Appendix, we determine the upper limit of the stabilizer Rényi entropy Leone et al. (2022) for generic two qubit systems with vanishing polarization. Following Liu et al. (2025b), we parametrize pure two-qubit states as
@@
It is easy to check that when (87) is satisfied, (85) is automatically satisfied. The maximal value of $M_{2}$ for systems with $\vec{B}_{1}=\vec{B}_{2}=0$ is therefore obtained by minimizing (16) of Liu et al. (2025b) with the constraints (87). There are many degenerate minima. For example, we find
@@
Minimizing (16) of Liu et al. (2025b) under this additional constraint, we find the extremal condition $\cos(4\phi_{1})=-1$, or equivalently,